# Beyond Community Notes: A Framework for Understanding and Building Crowdsourced Context Systems for Social Media

### v2

@@ Abstract.
Social media platforms are increasingly developing features that display crowdsourced context alongside posts, modeled after X’s Community Notes. These systems, which we term Crowdsourced Context Systems (CCS), have the potential to reshape our information ecosystem as major platforms embrace them as alternatives to top-down fact-checking. To deeply understand the features and implications of such systems, we perform a systematic literature review of existing CCS research and analyze several real-world CCS implementations. Based on our analysis, we develop a framework with two distinct components. First, we present a theoretical model to help conceptualize and define CCS. Second, we identify a design space encompassing six key aspects of CCS: participation, inputs, curation, presentation, platform treatment, and transparency. We discuss key normative implications of different CCS design and implementation choices. Our paper integrates these theoretical, design, and ethical perspectives to establish a foundation for future human-centered research on Crowdsourced Context Systems.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
Refer to caption Refer to caption :
Figure 1. Screenshots of a Community Note displayed on Instagram (left) and X (right).
@@ 1. Introduction
Social media platforms play an important role as sites of entertainment (herderDigitalJunkfoodSocial2024), news consumption (chouakiWhatNewsPeople2024a), and interpersonal connection (preeceEmpathicCommunitiesBalancing1999). At the same time, these platforms must contend with problems stemming from harmful user-uploaded content, which can take the form of abuse, spam, or misinformation (scheuermanFrameworkSeverityHarmful2021). Specifically, platforms have tried various forms of content moderation (grimmelmann2015virtues; lampeSlashdotBurnDistributed2004; jhaverHumanMachineCollaborationContent2019) to deal with misinformation, including centralized fact-checking efforts to label or remove it (gillespieCustodiansInternetPlatforms2018), as well as distributed approaches where users are empowered to tackle these challenges themselves (seeringReconsideringSelfModerationRole2020). While these approaches have had some success, they have also evoked backlash and accusations of bias from both users and lawmakers, particularly on the political right (jiangReasoningPoliticalBias2020; johnsonFactsAllegationsPolitical2023). Recently, a new approach has emerged: beginning with X/Twitter, major social media platforms are increasingly introducing systems that crowdsource “helpful context” to display alongside potentially problematic posts (IntroducingBirdwatchCommunitybaseda) (e.g. Figure 1). These systems, which we term Crowdsourced Context Systems (CCS), have seen growing adoption and impact, with some of the world’s largest platforms, including Meta (Facebook, Instagram, and Threads), TikTok, and YouTube all recently launching their own implementations. Meta has gone one step further, announcing that their CCS will replace their fact-checking program in the U.S. (heatheraMoreSpeechFewer2025).
@@
As CCS adoption grows, these systems have the potential to reshape our information ecosystem. Early studies into such systems have produced promising results, finding that CCS can increase user informedness and decrease the spread of misinformation (wojcikBirdwatchCrowdWisdom2022b; rollout21; diffusion21), although it is still unclear whether CCS can effectively replace centralized fact-checking, as some platforms have done or may seek to do (heatheraMoreSpeechFewer2025). As CCS develop, novel implementation questions also come to the fore, including how to design these systems most effectively—for example, to provide reliable context quickly enough to prevent the spread of the most viral content (rollout21; chuaiCommunitybasedFactcheckingReduces2024a; renaultCollaborativelyAddingContext2024a). At this critical moment of change in the information ecosystem, it is important to understand how CCS function, how we should evaluate them, and how these systems may be better designed.
@@
This paper makes two contributions towards the goal of better understanding CCS. Our first contribution is a systematic literature review of existing research studying CCS implementations. Our review finds significant research into CCS contributor behavior and the impact of displaying notes. However, the review highlights the fact that certain aspects of these systems are underexplored by existing research. We thus set out to identify additional facets of these systems that may affect their impact. Our second contribution is an analysis and comparison of real-world implementations of CCS by several major social media platforms, in order to create a framework to guide future HCI research on these novel and important systems. Our framework is composed of two parts. First, we offer a theoretical model of a CCS that identifies these systems’ core characteristics and what sets them apart from prior crowdsourced systems. Our model posits that while CCS use familiar moderation tools like crowdsourced information and context labels, they are unique in how they are natively integrated into platforms with the explicit goal of facilitating content interpretation. Second, we outline a design space (macleanQuestionsOptionsCriteria1996) for CCS in order to taxonomize the important design decisions made in their implementations. Design spaces are a conceptual tool used in HCI research that can guide work on complex socioetechnical systems by structuring the space of design possibilities (zhangFormFromDesignSpace2024a; leeDesignSpaceIntelligent2024; cardDesignSpaceInput1990). We group these design decisions into six high-level aspects of CCS that combine and interact to shape their impact: participation, inputs, curation, platform treatment, and transparency. Based on our framework, we discuss key normative implications of these systems in order to make sense of the ethical stakes involved in their design. As systems with the power to shape the information ecosystem, we emphasize three interconnected concerns: user informedness, distribution of power, and fairness. We situate CCS within the broader conversation about content moderation and establish a foundation for future human-centered research on Crowdsourced Context Systems.
@@ 2. Background
In January 2021, Twitter (now X) launched a U.S. pilot for the first large-scale CCS, which they called Birdwatch, presenting it as “a community-based approach to misinformation” (IntroducingBirdwatchCommunitybaseda). This system allowed a subset of users to suggest notes to “provide informative context” about Tweets (posts) on the platform. The same subset of users then rated these suggested notes and these ratings were used as the inputs to a note curation algorithm that aimed to select notes that would be helpful to a broad audience. In this pilot, these notes were only visible on a separate site, but Twitter announced their plan “to make notes visible directly on Tweets for the global Twitter audience, when there is consensus from a broad and diverse set of contributors.”
In March 2022, Twitter expanded the pilot, announcing that now “a small (and randomized) group of people on Twitter in the US will see Birdwatch notes directly on some Tweets” (BuildingBetterBirdwatch). Twitter was purchased by Elon Musk in October 2022, renamed X, and soon Birdwatch was renamed Community Notes. In December 2022, the pilot phase ended and users all over the world began seeing selected notes directly alongside Tweets (BeginningTodayCommunity2022). Since then, the system has continued to evolve. A key feature of X’s Community Notes is its transparency: the note curation code is open source [^1] https://github.com/twitter/communitynotes/tree/main, as are the crowdsourced notes and ratings data collected by the system [^2] https://x.com/i/communitynotes/download-data.
In June 2024, YouTube announced a pilot for a similar system, which they called Notes (TestingNewWays). In January 2025, Meta announced a pilot for their own implementation, also called Community Notes, which would launch on Facebook, Instagram, and Threads (heatheraMoreSpeechFewer2025). Finally, TikTok announced their own pilot in April 2025, which they called Footnotes (TestingNewFeature). According to the platforms, all of these new systems are based on X’s open source implementation of the note curation algorithm. Unlike X’s implementation, though, none of the new systems are open source or open data. In fact, while all of these systems serve similar purposes, they differ from each other in crucial ways due to unique design decisions made by each platform. Since these systems are increasingly prominent features of some of the most important social media platforms, it is critical to define, understand, and reason about them so the HCI community can better inform their implementation and evaluate their impact.
@@ 3. Related Work
While CCS are a relatively new innovation, these systems relate to and touch on multiple bodies of work from the HCI literature, including: research on content moderation in public discourse systems, research on fact-checking infrastructures, and research on crowdsourced responses to misinformation. We describe these areas in general terms here; later in the paper (Section 6) we provide a detailed review of the literature that is examining specific implementations of CCS.
@@
In a democracy that values free speech, public participation in civil discussion is considered a social good (smallPolisEscalarDeliberacion2021). One classic concern of HCI research has thus been how to design systems that can support public discussion among users with diverse, possibly conflicting, perspectives (kripleanSupportingReflectivePublic2012; wohnHowHandleOnline2017). CCS are an example of such a system. Like past public discussion systems, CCS provide a form of “content moderation,” which can be used to encourage certain types of participation and discourage others (grimmelmann2015virtues). Content moderation can take different approaches: it can be “top-down,” where platforms make moderation decisions (gillespieCustodiansInternetPlatforms2018; gillespieNotRecommendReduction2022), or “bottom-up,” where users do (seeringReconsideringSelfModerationRole2020; lloyd25; lloydAIRulesCharacterizing2025; fieslerRedditRulesCharacterizing2018). Additionally, moderation can be human-powered, or use varying degrees of automation to keep up with large volumes of content (vaccaroContestabilityContentModeration2021; gorwa20). Community Notes, X’s CCS implementation, incorporates all of these approaches: it uses a centralized algorithm to aggregate moderation feedback that is sourced from users. Community Notes is an example of a bridging system, or an automated system designed to, “increase mutual understanding and trust across divides, creating space for productive conflict, deliberation, or cooperation” (ovadyaBridgingSystemsOpen2023). Indeed, according to its creators, Community Notes was designed to identify crowdsourced content (notes) “that both inform understanding (decrease propensity to agree with a potentially misleading claim) and are seen as helpful by a diverse population of users” (wojcikBirdwatchCrowdWisdom2022b). CCS thus reflect a long-standing HCI concern with designing systems that support civil discussion through content moderation, adapting these ideas to meet the scale and speed of major social media platforms.
@@
One moderation approach frequently studied in HCI research is the use of top-down fact-checking to deal with misinformation (walterFactCheckingMetaAnalysisWhat2020; youngFactCheckingEffectivenessFunction2018). Emerging from journalistic practice, top-down fact-checking approaches empower certain individuals or organizations to verify public claims and publish determinations about their accuracy. While these techniques can have positive effects, they are time-consuming and traditionally require a large amount of professional labor (hassanDetectingCheckworthyFactual2015; junejaHumanTechnologicalInfrastructures2022), and thus may struggle to keep up with the speed at which content is published on the internet (micallefTrueFalseStudying2022). As a result, researchers have explored automated ways to scale the fact-checking process (warrenShowMeWork2025; guoSurveyAutomatedFactChecking2022; hassanQuestAutomateFactChecking). Beyond research into fact-check production, studies have also investigated the effects that fact-checks have on users who encounter them (eckerPsychologicalDriversMisinformation2022). This work has explored both the most effective ways to present fact-checks in order to shift user belief (claytonRealSolutionsFake2020), as well as the psychological characteristics that make some users reluctant to engage with them (tanakaWhoDoesNot2023). In addition to issues of efficacy, and the large cost of running such fact-check programs, social media platforms have struggled with allegations and perceptions that their fact-checks are biased against certain political ideologies (cotterFactCheckingCrisisCOVID192022; jiangReasoningPoliticalBias2020).
@@
CCS can be seen as an evolution in content moderation that aims to address the limitations of top-down fact-checking by empowering platform users. In doing so, these new systems build on a long history of HCI research that has explored ways of harnessing the distributed wisdom of a crowd to produce accurate information in networked settings (pennycookFightingMisinformationSocial2019; heSurveyRoleCrowds2025). This research has shown that crowds can perform various tasks, such as sourcing on-the-ground information (daileyJournalistsCrowdsourcerersResponding2014; agapieCrowdsourcingFieldCase2015), or fact-checking political claims (kripleanIntegratingOndemandFactchecking2014). Studies have also found that despite being cheaper and faster than professional fact-checking, crowd-based fact-checking can be similarly accurate and effective in certain circumstances  (allenScalingFactcheckingUsing2021; martelCrowdsCanEffectively2024a; zhaoInsightsComparativeStudy2023). At the same time, crowd-based fact-checking is not a silver bullet, and other studies have shown that the composition of the crowd can affect performance (godelModeratingMobEvaluating2021), and that users may respond negatively to “social correction” that comes from the crowd (mosleh21). Crowds that are representative of the general population may be biased and cannot accurately perform every task (drawsEffectsCrowdWorker2022). Researchers have studied techniques to mitigate these biases (spiekerDiversePerspectivesCan2023) and have explored other ways that crowds can play a useful role in moderating misleading information, such as via digital juries (fanDigitalJuriesCivicsOriented2020; zhaoEffectsNonEffectsSocial2025), or via distributed moderation with up/down votes (lampeSlashdotBurnDistributed2004; epsteinWillCrowdGame2020). CCS share much in common with previously explored crowdsourced fact-checking systems, but they have a novel set of processes and structures that lead us to define them as a new class of system. Informed by these past bodies of research, we next define what makes these new systems unique.
Figure 2. Model of a Crowdsourced Context System. Users create posts, context contributors create notes about those posts, and ratings contributors create ratings for those notes. Based on these ratings, the system selects “helpful” notes, which are displayed alongside the post for users to see.:
@@ 4. Generalized Model of Crowdsourced Context Systems for Social Media
We coin the term Crowdsourced Context System to describe these social media systems, a novel and important addition to the content moderation toolbox. Such systems have several core characteristics. First, CCS rely on contributions crowdsourced from users of the platform. Second, these systems are designed to specifically elicit context from the crowd that is meant to help users interpret the original post (as opposed to other types of responses—for example, the most humorous reply). Crowdsourced context can facilitate interpretation by providing additional information that helps readers assess and interpret the veracity, potential motivations, or broader social or political context of the original post. Third, these systems supplement user content by displaying additional text, or counterspeech (pengRescuingCounterspeechBridgingBased2024a), alongside the original posts. Finally, these systems are natively integrated into social media platforms. Unlike organic crowd behavior that emerges among users, these systems are built and operated by the platforms, giving platform designers the power to influence how the systems function.
Figure 2 shows a generalized model of a CCS. Such systems are composed of several core entities, which we denote with italics through the rest of the paper. Users create posts on the platform. Contributors then provide context, in the form of text or media snippets called notes, which are associated with specific posts. At first these notes are only visible to other contributors, who assess them and produce ratings of their helpfulness. The platform uses these ratings to identify “helpful” notes, which are then published and displayed alongside the associated post everywhere that it appears on the platform, as seen in Figure 1. In implementing this general model, platforms make many individual design decisions, which we will address later.

### v3

@@ Abstract.
Social media platforms are increasingly adopting features that display crowdsourced context alongside posts, a technique pioneered by X’s Community Notes. These systems—which we term Crowdsourced Context Systems (CCS)—have the potential to reshape the information ecosystem as major platforms embrace them as alternatives to professional fact-checking. To understand the features and implications of these systems, we conduct a systematic literature review of existing CCS research (n=56) and analyze real-world CCS implementations. Based on our analysis, we develop a framework with two components. First, we present a theoretical model to conceptualize and define CCS. Second, we identify a design space encompassing six aspects: participation, inputs, curation, presentation, platform treatment, and transparency. We also surface normative implications of different CCS design and implementation choices. Our work integrates theoretical, design, and ethical perspectives to establish a foundation for future human-centered research on Crowdsourced Context Systems.
Crowdsourced Context Systems, Community Notes, Social Media, Design Framework
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791889isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Collaborative and social computing systems and toolsccs: Human-centered computing Empirical studies in collaborative and social computing
Refer to caption Refer to caption Two side-by-side screenshots showing Community Notes displayed on different platforms. The left image shows Instagram with a post containing a photo of two people walking, with a Community Note appearing below providing the additional context that the image is AI. The right image shows X (formerly Twitter) with a post about Mars and a Community Note displayed underneath adding the additional context that the post is overstating the claims.
Figure 1. Screenshots of crowdsourced notes displayed on Instagram (left) and X (right).
@@ 1. Introduction
Social media platforms are important sites of entertainment (Herder and Staring, 2024), news consumption (Chouaki et al., 2024), and interpersonal connection (Preece, 1999). However, these platforms must contend with harmful user-uploaded content, including abuse, spam, and misinformation (Scheuerman et al., 2021). To address misinformation, platforms have tried various content moderation approaches (Grimmelmann, 2015; Lampe and Resnick, 2004; Jhaver et al., 2019), from using professional fact-checkers to label or remove false content (Gillespie, 2018) to empowering user communities to moderate themselves (Seering, 2020). While professional fact-checking has been somewhat successful, it has also evoked backlash and accusations of bias from users and lawmakers, particularly on the political right (Jiang et al., 2020; Johnson, 2023). Recently, a new approach has emerged: beginning with X/Twitter (Coleman, 2021), major platforms are increasingly introducing systems that crowdsource “helpful context” from users to display alongside potentially problematic posts (see examples in Figure 1). These systems, which we term Crowdsourced Context Systems (CCS), have seen growing adoption, with some of the world’s largest platforms—including Meta (Facebook, Instagram, and Threads), TikTok, and YouTube—all recently launching implementations. Meta demonstrated the importance of these systems when they announced in January 2025 that their CCS would replace their U.S. fact-checking program (Kaplan, 2025).
@@
As CCS adoption grows, these systems have the potential to reshape the information ecosystem. Early studies have produced promising results, finding that CCS can increase user informedness and decrease misinformation spread (Wojcik et al., 2022; Chuai et al., 2024b; Drolsbach and Pröllochs, 2023), though it remains unclear whether CCS can effectively replace professional fact-checking, as Meta is attempting (Kaplan, 2025). As CCS develop, novel implementation questions come to the fore, including how to best design these systems to achieve their stated or implicit goals—for example, to provide reliable context quickly enough to prevent the spread of viral misleading content (Chuai et al., 2024b, a; Renault et al., 2024). At this critical juncture in our information ecosystem’s development, it is important to understand how CCS function, how to evaluate them, and how to improve their design.
@@
This paper makes two contributions toward better understanding CCS. Our first contribution is a systematic literature review of CCS research. Our review finds significant work on CCS contributor behavior and the impact of displaying context notes alongside problematic posts. However, the review also reveals that certain aspects of these systems remain underexplored. Our second contribution addresses this gap: an analysis and comparison of real-world CCS implementations across major social media platforms. From this analysis, we create a framework to guide future HCI research on CCS. Our framework has two parts. First, we offer a theoretical model that identifies CCS core characteristics and what distinguishes them from prior crowdsourced systems. Our model posits that while CCS use familiar moderation tools like crowdsourced information and context labels, they are unique in being natively integrated into platforms with the explicit goal of facilitating content interpretation. Second, we outline a design space (MacLean et al., 1996) for CCS to taxonomize important implementation decisions. Design spaces are a conceptual HCI tool that guide work on complex socioetechnical systems by structuring design possibilities (Zhang et al., 2024; Lee et al., 2024; Card et al., 1990). We group these decisions into six high-level aspects that combine and interact to shape CCS impact: participation, inputs, curation, presentation, platform treatment, and transparency.
Based on our framework, the Discussion section highlights key normative implications of CCS in order to make sense of the ethical stakes involved in their design. As these systems aim to shape the information ecosystem, we emphasize three interconnected normative concerns: user informedness, distribution of power, and fairness. Additionally, we discuss how CCS fit within the broader conversation about content moderation and establish a foundation for future human-centered research on Crowdsourced Context Systems.
@@ 2. Background
In January 2021, Twitter (now X) launched a U.S. pilot for the first large-scale CCS, which they called Birdwatch, announcing it as “a community-based approach to misinformation” (Coleman, 2021). This system allowed a subset of users to suggest notes to “provide informative context” about Tweets (posts) on the platform. The same users then rated these suggested notes, and the ratings were used as input for a note curation algorithm that aimed to select notes that would be helpful to a broad audience. In this pilot, these notes were only visible on a separate site, but Twitter announced a plan: “to make notes visible directly on Tweets for the global Twitter audience, when there is consensus from a broad and diverse set of contributors.”
In March 2022, Twitter expanded the pilot, announcing that: “a small (and randomized) group of people on Twitter in the US will see Birdwatch notes directly on some Tweets” (Coleman, 2022). Twitter was purchased by Elon Musk in October 2022, renamed X, and soon Birdwatch was renamed Community Notes. In December 2022, the pilot phase ended and users all over the world began seeing selected notes directly alongside Tweets (@CommunityNotes, 2022). Since then, the system has continued to evolve. A key feature of X’s Community Notes is its transparency: the note curation code is open source [^1] https://github.com/twitter/communitynotes/tree/main, as are the crowdsourced notes and ratings data collected by the system [^2] https://x.com/i/communitynotes/download-data.
In June 2024, YouTube announced a pilot for a similar system, which they called Notes (Team, 2024). In January 2025, Meta announced a pilot for their own implementation, also called Community Notes, which would launch on Facebook, Instagram, and Threads (Kaplan, 2025). Finally, TikTok announced their own pilot in April 2025, which they called Footnotes (Presser, 2025). At the time of writing, all of these systems have launched, at least as limited pilots. According to the platforms, all of these systems are based on X’s open source note curation algorithm. Unlike X’s implementation, though, none of the new systems make their code or data public. In fact, while they serve similar purposes, these systems differ from each other in crucial ways due to platform design decisions. As CCS proliferate on major platforms, the HCI community must define, understand, and reason about them to evaluate their impact and inform better implementations.
@@ 3. Related Work
While CCS are a relatively new innovation, these systems relate to and touch on multiple bodies of work from the HCI literature, including: research on content moderation in public discourse systems, research on fact-checking infrastructures, and research on crowdsourced responses to misinformation. We describe these areas generally here; later in Section 6 we provide a detailed review of research that examines specific implementations of CCS.
@@
In a democracy that values free speech, public participation in civil discussion is considered a social good (Small, 2021). One classic concern of HCI research has thus been how to design systems that can support public discussion among users with diverse, possibly conflicting, perspectives (Kriplean et al., 2012; Wohn et al., 2017). CCS are an example of such a system. Like past public discussion systems, CCS provide a form of content moderation that encourages certain types of participation and discourages others (Grimmelmann, 2015). Content moderation can take different approaches: it can be “top-down,” where platforms make moderation decisions (Gillespie, 2018, 2022), or “bottom-up,” where users do (Seering, 2020; Lloyd et al., 2025b, a; Fiesler et al., 2018). Additionally, moderation can be human-powered, or use varying degrees of automation to keep up with large volumes of content (Vaccaro et al., 2021; Gorwa et al., 2020). Community Notes, X’s CCS implementation, incorporates all of these approaches: it uses a centralized algorithm to aggregate moderation feedback that is sourced from users and, as of late 2025, user-created AI note writers [^3] https://x.com/CommunityNotes/status/1940132205486915917. X’s Community Notes could be seen as an example of a bridging system, or an automated system designed to “increase mutual understanding and trust across divides, creating space for productive conflict, deliberation, or cooperation” (Ovadya and Thorburn, 2023). Indeed, according to its creators, Community Notes was designed to identify crowdsourced content (notes) “that both inform understanding (decrease propensity to agree with a potentially misleading claim) and are seen as helpful by a diverse population of users” (Wojcik et al., 2022). CCS thus reflect a long-standing HCI concern with designing systems that support civil discussion through content moderation, adapting these ideas to meet the scale and speed of major social media platforms.
@@
One moderation approach frequently studied in HCI research is the use of professional fact-checking to deal with misinformation (Walter et al., 2020; Young et al., 2018). Emerging from journalistic practice, professional fact-checking empowers certain individuals or organizations to verify public claims and publish determinations about their accuracy. While these techniques can have positive effects, they are time-consuming and require a large amount of professional labor (Hassan et al., 2015b; Juneja and Mitra, 2022), and thus may struggle to keep up with the speed at which content is published on the internet (Micallef et al., 2022). As a result, researchers have explored automated ways to scale fact-checking (Warren et al., 2025; Guo et al., 2022; Hassan et al., 2015a). Beyond research into fact-check production, studies have also investigated the effects that fact-checks have on users who encounter them (Ecker et al., 2022). This work has explored both the most effective ways to present fact-checks in order to shift user belief (Clayton et al., 2020), as well as the psychological characteristics that make some users reluctant to engage with them (Tanaka et al., 2023). In addition to issues of efficacy, and the large cost of running such fact-check programs, social media platforms have struggled with allegations and perceptions that their fact-checks are biased against certain political ideologies (Cotter et al., 2022; Jiang et al., 2020).
@@
CCS can be seen as an evolution in content moderation that aims to address the limitations of professional fact-checking by enlisting platform users. In doing so, these new systems build on a long history of HCI research exploring ways of harnessing the distributed wisdom of a crowd to produce accurate information in networked settings (Pennycook and Rand, 2019; He et al., 2025). This research has shown that crowds can effectively perform various tasks, such as sourcing on-the-ground information (Dailey and Starbird, 2014; Agapie et al., 2015), or fact-checking political claims (Kriplean et al., 2014). Studies have also found that despite being cheaper and faster than professional fact-checking, crowd-based fact-checking can be similarly accurate in certain circumstances  (Allen et al., 2021; Martel et al., 2024; Zhao and Naaman, 2023). However, crowd-based fact-checking is not a silver bullet. Users may respond negatively to “social correction” from peers (Mosleh et al., 2021), the composition of the crowd affects its ability to identify false claims (Godel et al., 2021), and crowds representative of the general population can exhibit biases that limit their accuracy (Draws et al., 2022). Researchers have studied techniques to mitigate these fact-checking biases (Thebault-Spieker et al., 2023) and have explored other ways that crowds can moderate misleading information, such as via digital juries (Fan and Zhang, 2020; Zhao and Hobbs, 2025), or via distributed moderation with up/down votes (Lampe and Resnick, 2004; Epstein et al., 2020). CCS share much in common with previously explored crowdsourced fact-checking systems, but they have a novel set of processes and structures that lead us to define them as a new class of system. Informed by these past bodies of research, we next define what makes these new systems unique.
@@ 4. Generalized Model of Crowdsourced Context Systems for Social Media
Figure 2. Model of a Crowdsourced Context System. Users create posts, context contributors create notes about those posts, and ratings contributors create ratings for those notes. Based on these ratings, the system selects “helpful” notes, which are displayed alongside the post for users to see.
A flowchart diagram illustrating the Crowdsourced Context System model with 5 numbered steps: (1) User creates post, (2) Context Contributors create Notes for post, (3) Rating Contributors create Ratings for Notes, (4) System selects ”Helpful” Notes based on Ratings, and (5) Platform displays selected Notes on post to provide context. The diagram shows the flow between users, the crowdsourced context system, and the final display.
We coin the term Crowdsourced Context System to describe these social media systems, a novel and important addition to the content moderation toolbox. Such systems have several core characteristics. First, CCS rely on contributions crowdsourced from users of the platform. Second, these systems are designed to specifically elicit and identify context from the crowd that can help users interpret the original post. Crowdsourced context can facilitate interpretation by providing additional information that helps readers assess and interpret the veracity, potential motivations, or broader social or political context of the original post. This explicit focus on facilitating interpretation differentiates CCS from earlier crowdsourced systems which may have more general aims, such as traditional social media comment sections that emphasize the most engaging responses. Third, these systems do not remove user content, but supplement it by displaying additional text, or counterspeech (Peng and Grimmelmann, 2024), alongside the original posts. Finally, these systems are natively integrated into social media platforms. Unlike crowd behavior that emerges organically among users, these systems are built and operated by the platforms themselves, giving platform designers the power to influence how the systems function.
Figure 2 shows a generalized model of a CCS. Such systems are composed of several core entities. Users create posts on the platform. Contributors then provide context, in the form of text or media snippets called notes, which are associated with specific posts. At first these notes are only visible to other contributors, who assess them and produce ratings of their helpfulness. The platform uses these ratings to identify “helpful” notes, which are then published and displayed alongside the associated post everywhere that it appears on the platform, as seen in Figure 1. In implementing this general model, platforms make many individual design decisions, which we highlight and outline below.